# Learning Theory for Kernel Bilevel Optimization

### v1

@@ D.1 Point-wise estimates
$\begin{split}{\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}&\leq\underbrace{{\left\lVert\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert}}_{\partial_{\omega}\delta_{\omega}^{out}}+\underbrace{{\left\lVert\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rVert}}_{\partial_{\omega}E_{\omega}^{out}}\\ +&{\left\lVert\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}{\left(\underbrace{{\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert}_{\mathcal{H}}}_{\partial_{h}\delta_{\omega}^{out}}+\underbrace{{\left\lVert\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rVert}_{\mathcal{H}}}_{\partial_{h}E_{\omega}^{out}}\right)}\\ +&{\left\lVert R(h^{\star}_{\omega})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}{\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})\right\rVert}_{\mathcal{H}}.\end{split}$ (17)
Next, we provide upper-bounds on ${\left\lVert R(h_{\omega}^{\star})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$ and ${\left\lVert\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$ in terms of derivatives of $L_{in}$ and $\widehat{L}_{in}$.
Upper-bounds on ${\left\lVert R(h_{\omega}^{\star})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$ and ${\left\lVert\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$. By application of Propositions A.2 and A.3, we deduce that $\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})$, $\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})$, $\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})$ and $\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})$ are all bounded operators. Moreover, since $L_{in}$ and $\widehat{L}_{in}$ are $\lambda$-strongly convex in their second argument by Proposition A.3, it follows that $\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})\geq\lambda\operatorname{Id}_{\mathcal{H}}$ and $\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\geq\lambda\operatorname{Id}_{\mathcal{H}}$. We can therefore apply Lemma G.1 which yields the following inequalities:
@@
${\left\lVert R(h_{\omega}^{\star})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}\leq$ $\frac{1}{\lambda^{2}}{\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})\right\rVert}_{\operatorname{op}}{\left\lVert\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$
@@
${\left\lVert\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}\leq$ $\frac{1}{\lambda}{\left\lVert\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}.$
@@
${\left\lVert R(h_{\omega}^{\star})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}\leq$ $\frac{1}{\lambda^{2}}{\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})\right\rVert}_{\operatorname{op}}{\left(\underbrace{{\left\lVert\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert}_{\operatorname{op}}}_{\partial_{h}^{2}\delta_{\omega}^{in}}+\underbrace{{\left\lVert\partial_{h}^{2}\widehat{L}_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}}_{\partial_{h}^{2}E_{\omega}^{in}}\right)}$
@@
Final bound. We can now substitute the above bounds on ${\left\lVert R(h_{\omega}^{\star})-\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$ and ${\left\lVert\hat{R}(\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$ into Equation 17 to obtain the following upper-bound on the gradient error:
${\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}$ $\leq\partial_{\omega}\delta_{\omega}^{out}+\partial_{\omega}E_{\omega}^{out}+\frac{1}{\lambda}{\left\lVert\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}{\left(\partial_{h}\delta_{\omega}^{out}+\partial_{h}E_{\omega}^{out}\right)}$ (18)
$+$ ${\left\lVert\partial_{h}L_{out}(\omega,h_{\omega}^{\star})\right\rVert}_{\mathcal{H}}{\left(\frac{1}{\lambda^{2}}{\left\lVert\partial^{2}_{\omega,h}L_{in}(\omega,h_{\omega}^{\star})\right\rVert}_{\operatorname{op}}{\left(\partial_{h}^{2}\delta_{\omega}^{in}+\partial_{h}^{2}E_{\omega}^{in}\right)}+\frac{1}{\lambda}{\left(\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\partial_{\omega,h}^{2}E_{\omega}^{in}\right)}\right)}.$
Furthermore, by Proposition D.3, we have the following upper-bounds on the derivatives of $L_{in}$ and $L_{out}$:
@@
Incorporating the above bounds into Equation 18, we further get:
@@
By Proposition D.2, we can upper-bound the error terms $\partial_{\omega}E_{\omega}^{out},\partial_{h}E_{\omega}^{out}$ by $C_{out}{\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}$ and $\partial_{h}^{2}E_{\omega}^{in},\partial_{\omega,h}^{2}E_{\omega}^{in}$ by the difference $C_{in}{\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}$. Furthermore, since ${\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}\leq\frac{1}{\lambda}\partial_{h}\delta_{\omega}^{in}$ by Proposition D.1, we can further show that gradient error satisfies the desired bound:
${\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}\leq\partial_{\omega}\delta_{\omega}^{out}+\frac{C_{in}}{\lambda}\partial_{h}\delta_{\omega}^{out}+C_{out}{\left(\frac{C_{in}}{\lambda^{2}}\partial_{h}^{2}\delta_{\omega}^{in}+\frac{1}{\lambda}\partial_{\omega,h}^{2}\delta_{\omega}^{in}\right)}+\frac{C_{out}}{\lambda}{\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)}\partial_{h}\delta_{\omega}^{in}.$
D.2 Maximal inequalities
Proposition D.5 (Maximal inequalities for empirical processes).
Let $\Lambda$ be a positive constant. Under Assumptions (B), (A), (C) and (D), the following maximal inequalities hold for any $0<\lambda\leq\Lambda$:
$\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\delta_{\omega}^{out}\right]}$ $\leq{\sqrt{\frac{1}{\lambda^{2}m}}c(\Omega)\max(M_{out}{\operatorname{Lip}_{out}}\operatorname{diam}(\Omega),\Lambda{M_{out}^{2}})}$
@@ D.2 Maximal inequalities
where $c(\Omega)$ is a positive constant greater than 1 that depends only on $\Omega$ and $d$, while $\operatorname{Lip}_{out}$ and $M_{out}$ are positive constants defined in Propositions C.3 and C.4.
@@
We will apply the result of Proposition E.3 which provides maximal inequalities for real-valued empirical processes that are uniformly bounded and Lipschitz in their parameter. To this end, consider the parametric families:
@@
For any $0\leq l\leq d$, these real-valued functions are uniformly bounded by a positive constant $M_{out}$, thanks to Proposition C.3. Moreover, by Proposition C.4, the functions $\omega\mapsto\partial_{\omega_{l}}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$, and $\omega\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$ are all $\lambda^{-1}\operatorname{Lip}_{out}$-Lipschitz for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Hence, Proposition E.3 is applicable to each of these families, with $\mathbb{D}$ set to $\mathbb{Q}$ and $\mathcal{Z}$ set to $\mathcal{X}\times\mathcal{Y}$. We treat both $\delta_{\omega}^{out}$ and $\partial_{\omega}\delta_{\omega}^{out}$ separately.
A maximal inequality for $\delta_{\omega}^{out}$. For $l=0$, we readily apply Proposition E.3 with $p=1$ to get the following maximal inequality for $\delta_{\omega}^{out}$:
@@
$\leq$ $\sqrt{\frac{1}{\lambda^{2}m}}c(\Omega)\max({\operatorname{Lip}_{out}}\operatorname{diam}(\Omega),\Lambda{M_{out}^{2}}).$
A maximal inequality for $\partial_{\omega}\delta_{\omega}^{out}$. We now turn to $\partial_{\omega}\delta_{\omega}^{out}$, which involves vector-valued processes (as an error between the gradient and its estimate). While the maximal inequalities in Proposition E.3 hold for real-valued processes, we will first obtain maximal inequalities for each component appearing in $\partial_{\omega}\delta_{\omega}^{out}$ and then sum these to control $\partial_{\omega}\delta_{\omega}^{out}$. To this end, we first use the Cauchy-Schwarz inequality which implies that $\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})\right]}\leq\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]}^{\frac{1}{2}}$. Thus we only need to control $\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]}$. Simple calculations show that:
$\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\partial_{\omega}\delta_{\omega}^{out}\right]}^{2}$ $\leq\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]}$
@@
$\leq{\left(\sqrt{\frac{d}{\lambda^{2}m}}c(\Omega)\max(\operatorname{Lip}_{out}\operatorname{diam}(\Omega),\Lambda M_{out}^{{2}})\right)}^{2},$
where the last inequality follows by application of Proposition E.3 with $p=2$ to each term in the right-hand side of the first inequality for $1\leq l\leq d$. We get the desired bound on $\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\partial_{\omega}\delta_{\omega}^{out}\right]}$ by taking the square root of the above inequality. ∎
Proposition D.6 (Maximal inequalities for RKHS-valued empirical processes).
Let $\Lambda$ be a positive constant. Under Assumptions (B), (A), (C) and (D), the following maximal inequalities hold for any $0<\lambda\leq\Lambda$:
@@
where $c(\Omega)$ is a positive constant greater than $1$ that depends only on $\Omega$ and $d$, $\widetilde{L}_{out,1},\widetilde{L}_{in,1},\widetilde{L}_{in,2},\widetilde{M}_{out,1},\widetilde{M}_{in,1},\widetilde{M}_{in,2}$ are positive constants defined as:
@@
and $\operatorname{Lip}_{out},\operatorname{Lip}_{in},M_{out},M_{in}$ are positive constants given in Propositions C.3 and C.4.
@@
where $a\in\{1,2\}$, $s$ is an integer satisfying $20\leq s\leq d+$ 2, and $f_{s}(\omega,x,y)$ are real-valued functions given by:
$f_{0}:(\omega,x,y)\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y),\qquad f_{1}:(\omega,x,y)\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),\qquad f_{2}:(\omega,x,y)\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),$
$f_{2+l}:(\omega,x,y)\mapsto\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),\quad 1\leq l\leq d.$
For any $21\leq s\leq d+$ 2, the real-valued functions $f_{s}$ are uniformly bounded by a positive constant $M_{in}$ thanks to Proposition C.3. Moreover, since the kernel $K$ is bounded by $\kappa$ due to Assumption (A), it follows that all elements $t_{\omega}$ of $\mathcal{T}_{s,a}$ are uniformly bounded by $\widetilde{M}_{in,a}\coloneqq M_{in}^{2}\kappa^{a}$. Moreover, for $21\leq s\leq d+$ 2, the functions $\omega\mapsto f_{s}(\omega,x,y)$, are $\lambda^{-1}\operatorname{Lip}_{in}$-Lipschitz for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$, by Proposition C.4. Hence, it follows that the maps $\omega\mapsto t_{\omega}((x,y),(x^{\prime},y^{\prime}))$ are $\lambda^{-1}\widetilde{L}_{in,a}$-Lipschitz with $\widetilde{L}_{in,a}\coloneqq 2\operatorname{Lip}_{in}M_{in}\kappa^{a}$ for any $(x,y)$ and $(x^{\prime},y^{\prime})$ in $\mathcal{X}\times\mathcal{Y}$. Similarly, for $s=0$, we get the same properties, albeit, with different constants, i.e., the family $\mathcal{T}_{0,a}$ is uniformly bounded by a constant $\widetilde{M}_{out,a}\coloneqq M_{out}^{2}\kappa^{a}$ with $M_{out}$ introduced in Proposition C.3, and is $\lambda^{-1}\widetilde{L}_{out,a}$-Lipschitz in its parameter with $\widetilde{L}_{out,a}\coloneqq 2\operatorname{Lip}_{out}M_{out}\kappa^{a}$ where $\operatorname{Lip}_{out}$ is given in Proposition C.4. Hence, the maximal inequality in Proposition E.4 is applicable to each of these families with $\mathcal{Z}$ set to $\mathcal{X}\times\mathcal{Y}$, and $\mathbb{D}$ set either to $\mathbb{P}$ for $21\leq s\leq d+$ 2, or to $\mathbb{Q}$ for $s=0$. For conciseness, in all what follows, we will write $z=(x,y)$ and $z_{i}=(x_{i},y_{i})$ and $\tilde{z}_{j}=(\tilde{x}_{j},\tilde{y}_{j})$ for $1\leq i\leq n$ and $1\leq j\leq m$.
@@
$\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{out}\right]}^{2}$ $\leq\mathbb{E}_{\mathbb{Q}}{\left[\sup_{\omega\in\Omega}(\partial_{h}\delta_{\omega}^{out})^{2}\right]}$

### v3

@@ E.1 Point-wise estimates
$\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert$
@@
$\leq$ $\underbrace{\left\lVert\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert}_{\partial_{\omega}\delta_{\omega}^{out}}+\underbrace{\left\lVert\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rVert}_{\partial_{\omega}E_{\omega}^{out}}$
$+\left\lVert\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}\left(\underbrace{\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert_{\mathcal{H}}}_{\partial_{h}\delta_{\omega}^{out}}+\underbrace{\left\lVert\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rVert_{\mathcal{H}}}_{\partial_{h}E_{\omega}^{out}}\right)$
$+\left\lVert R(h^{\star}_{\omega})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})\right\rVert_{\mathcal{H}}.$ (16)
Next, we provide upper bounds on $\left\lVert R(h_{\omega}^{\star})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$ and $\left\lVert\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$ in terms of derivatives of $L_{in}$ and $\widehat{L}_{in}$.
Upper bounds on $\left\lVert R(h_{\omega}^{\star})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$ and $\left\lVert\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$. By application of Propositions˜B.2 and B.3, we deduce that $\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})$, $\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})$, $\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})$, and $\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})$ are all bounded operators. Moreover, since $L_{in}$ and $\widehat{L}_{in}$ are $\lambda$-strongly convex in their second argument by Proposition˜B.3, it follows that $\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})\geq\lambda\operatorname{Id}_{\mathcal{H}}$ and $\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\geq\lambda\operatorname{Id}_{\mathcal{H}}$. We can therefore apply Lemma˜H.1 which yields the following inequalities:
$\left\lVert R(h_{\omega}^{\star})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}\leq$ $\frac{1}{\lambda^{2}}\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})\right\rVert_{\operatorname{op}}\left\lVert\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$
@@
$\left\lVert\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}\leq$ $\frac{1}{\lambda}\left\lVert\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert_{\operatorname{op}}.$
@@
$\left\lVert R(h_{\omega}^{\star})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$
$\leq$ $\frac{1}{\lambda^{2}}\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h_{\omega}^{\star})\right\rVert_{\operatorname{op}}\left(\underbrace{\left\lVert\partial_{h}^{2}L_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert_{\operatorname{op}}}_{\partial_{h}^{2}\delta_{\omega}^{in}}+\underbrace{\left\lVert\partial_{h}^{2}\widehat{L}_{in}(\omega,h_{\omega}^{\star})-\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert_{\operatorname{op}}}_{\partial_{h}^{2}E_{\omega}^{in}}\right)$
@@
Final bound. We can now substitute the above bounds on $\left\lVert R(h_{\omega}^{\star})-\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$ and $\left\lVert\widehat{R}(\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$ into Section˜E.1 to obtain the following upper bound on the gradient error:
$\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert$
$\leq$ $\partial_{\omega}\delta_{\omega}^{out}+\partial_{\omega}E_{\omega}^{out}+\frac{1}{\lambda}\left\lVert\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert_{\operatorname{op}}\left(\partial_{h}\delta_{\omega}^{out}+\partial_{h}E_{\omega}^{out}\right)$
$+\left\lVert\partial_{h}L_{out}(\omega,h_{\omega}^{\star})\right\rVert_{\mathcal{H}}\left(\frac{1}{\lambda^{2}}\left\lVert\partial^{2}_{\omega,h}L_{in}(\omega,h_{\omega}^{\star})\right\rVert_{\operatorname{op}}\left(\partial_{h}^{2}\delta_{\omega}^{in}+\partial_{h}^{2}E_{\omega}^{in}\right)+\frac{1}{\lambda}\left(\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\partial_{\omega,h}^{2}E_{\omega}^{in}\right)\right).$ (17)
Furthermore, by Proposition˜E.3, we have the following upper bounds on the derivatives of $L_{in}$ and $L_{out}$:
@@
Incorporating the above bounds into Section˜E.1, we further get:
@@
By Proposition˜E.2, we can upper-bound the error terms $\partial_{\omega}E_{\omega}^{out}$ and $\partial_{h}E_{\omega}^{out}$ by $C_{out}\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}$, and $\partial_{h}^{2}E_{\omega}^{in}$ and $\partial_{\omega,h}^{2}E_{\omega}^{in}$ by $C_{in}\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}$. Furthermore, since $\left\lVert h_{\omega}^{\star}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}\leq\frac{1}{\lambda}\partial_{h}\delta_{\omega}^{in}$ by Proposition˜E.1, we can further show that the gradient error satisfies the desired bound:
$\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert\leq$ $\partial_{\omega}\delta_{\omega}^{out}+\frac{C_{in}}{\lambda}\partial_{h}\delta_{\omega}^{out}+C_{out}\left(\frac{C_{in}}{\lambda^{2}}\partial_{h}^{2}\delta_{\omega}^{in}+\frac{1}{\lambda}\partial_{\omega,h}^{2}\delta_{\omega}^{in}\right)$
$+\frac{C_{out}}{\lambda}\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)\partial_{h}\delta_{\omega}^{in}.$
E.2 Maximal inequalities
Proposition E.5 (Maximal inequalities for empirical processes).
Let $\Lambda$ be a positive constant. Under Assumptions˜(A), (C), (B), (D) and (E), the following maximal inequalities hold for any $0<\lambda\leq\Lambda$:
$\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\delta_{\omega}^{out}\right]$ $\leq{\sqrt{\frac{1}{\lambda^{2}m}}c(\Omega)\max(M_{out}{\operatorname{Lip}_{out}}\operatorname{diam}(\Omega),\Lambda{M_{out}^{2}})},$
@@ E.2 Maximal inequalities
where $c(\Omega)$ is a positive constant greater than 1 that depends only on $\Omega$ and $d$, while $\operatorname{Lip}_{out}$ and $M_{out}$ are positive constants defined in Propositions˜D.3 and D.4.
@@
We will apply the result of Proposition˜F.3 which provides maximal inequalities for real-valued empirical processes that are uniformly bounded and Lipschitz in their parameter. To this end, consider the parametric families:
@@
For any $0\leq l\leq d$, these real-valued functions are uniformly bounded by a positive constant $M_{out}$, thanks to Proposition˜D.3. Moreover, by Proposition˜D.4, the functions $\omega\mapsto\partial_{\omega_{l}}\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$ and $\omega\mapsto\ell_{out}(\omega,h_{\omega}^{\star}(x),y)$ are all $\lambda^{-1}\operatorname{Lip}_{out}$-Lipschitz for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Hence, Proposition˜F.3 is applicable to each of these families, with $\mathbb{D}$ set to $\mathbb{Q}$ and $\mathcal{Z}$ set to $\mathcal{X}\times\mathcal{Y}$. We treat both $\delta_{\omega}^{out}$ and $\partial_{\omega}\delta_{\omega}^{out}$ separately.
A maximal inequality for $\delta_{\omega}^{out}$. For $l=0$, we readily apply Proposition˜F.3 with $p=1$ to get the following maximal inequality for $\delta_{\omega}^{out}$:
@@
$\leq$ $\sqrt{\frac{1}{\lambda^{2}m}}c(\Omega)\max(M_{out}{\operatorname{Lip}_{out}}\operatorname{diam}(\Omega),\Lambda{M_{out}^{2}}).$
A maximal inequality for $\partial_{\omega}\delta_{\omega}^{out}$. We now turn to $\partial_{\omega}\delta_{\omega}^{out}$, which involves vector-valued processes (as an error between the gradient and its estimate). While the maximal inequalities in Proposition˜F.3 hold for real-valued processes, we will first obtain maximal inequalities for each component appearing in $\partial_{\omega}\delta_{\omega}^{out}$ and then sum these to control $\partial_{\omega}\delta_{\omega}^{out}$. To this end, we first use the Cauchy-Schwarz inequality which implies that $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\partial_{\omega}\delta_{\omega}^{out}\right]\leq\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]^{\frac{1}{2}}$. Thus we only need to control $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]$. Simple calculations show that:
$\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\partial_{\omega}\delta_{\omega}^{out}\right]^{2}$
$\leq$ $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}(\partial_{\omega}\delta_{\omega}^{out})^{2}\right]$
@@
$\leq$ $\left(\sqrt{\frac{d}{\lambda^{2}m}}c(\Omega)\max(M_{out}\operatorname{Lip}_{out}\operatorname{diam}(\Omega),\Lambda M_{out}^{{2}})\right)^{2},$
where the last inequality follows by application of Proposition˜F.3 with $p=2$ to each term in the right-hand side of the first inequality for $1\leq l\leq d$. We get the desired bound on $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\partial_{\omega}\delta_{\omega}^{out}\right]$ by taking the square root of the above inequality. ∎
Proposition E.6 (Maximal inequalities for RKHS-valued empirical processes).
Let $\Lambda$ be a positive constant. Under Assumptions˜(A), (C), (B), (D) and (E), the following maximal inequalities hold for any $0<\lambda\leq\Lambda$:
@@
where $c(\Omega)$ is a positive constant greater than $1$ that depends only on $\Omega$ and $d$, $\widetilde{L}_{out,1},\widetilde{L}_{in,1},\widetilde{L}_{in,2},\widetilde{M}_{out,1},\widetilde{M}_{in,1}$, and $\widetilde{M}_{in,2}$ are positive constants defined as:
@@
and $\operatorname{Lip}_{out},\operatorname{Lip}_{in},M_{out}$, and $M_{in}$ are positive constants given in Propositions˜D.3 and D.4.
@@
where $a\in\{1,2\}$, $s$ is an integer satisfying $0\leq s\leq d+2$, and $f_{s}(\omega,x,y)$ are real-valued functions given by:
$f_{0}:(\omega,x,y)\mapsto\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y),\qquad f_{1}:(\omega,x,y)\mapsto\partial_{v}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),$
$f_{2}:(\omega,x,y)\mapsto\partial_{v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),\qquad f_{2+l}:(\omega,x,y)\mapsto\partial_{\omega_{l},v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y),\quad 1\leq l\leq d.$
For any $1\leq s\leq d+2$, the real-valued functions $f_{s}$ are uniformly bounded by a positive constant $M_{in}$ thanks to Proposition˜D.3. Moreover, since the kernel $K$ is bounded by $\kappa$ due to Assumption˜(B), it follows that all elements $t_{\omega}$ of $\mathcal{T}_{s,a}$ are uniformly bounded by $\widetilde{M}_{in,a}\coloneqq M_{in}^{2}\kappa^{a}$. Moreover, for $1\leq s\leq d+2$, the functions $\omega\mapsto f_{s}(\omega,x,y)$ are $\lambda^{-1}\operatorname{Lip}_{in}$-Lipschitz for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ by Proposition˜D.4. Hence, it follows that the map $\omega\mapsto t_{\omega}((x,y),(x^{\prime},y^{\prime}))$ is $\lambda^{-1}\widetilde{L}_{in,a}$-Lipschitz with $\widetilde{L}_{in,a}\coloneqq 2\operatorname{Lip}_{in}M_{in}\kappa^{a}$ for any $(x,y)$ and $(x^{\prime},y^{\prime})$ in $\mathcal{X}\times\mathcal{Y}$. Similarly, for $s=0$, we get the same properties, albeit, with different constants, i.e., the family $\mathcal{T}_{0,a}$ is uniformly bounded by a constant $\widetilde{M}_{out,a}\coloneqq M_{out}^{2}\kappa^{a}$ with $M_{out}$ introduced in Proposition˜D.3, and is $\lambda^{-1}\widetilde{L}_{out,a}$-Lipschitz in its parameter with $\widetilde{L}_{out,a}\coloneqq 2\operatorname{Lip}_{out}M_{out}\kappa^{a}$ where $\operatorname{Lip}_{out}$ is given in Proposition˜D.4. Hence, the maximal inequality in Proposition˜F.4 is applicable to each of these families with $\mathcal{Z}$ set to $\mathcal{X}\times\mathcal{Y}$, and $\mathbb{D}$ set either to $\mathbb{P}$ for $1\leq s\leq d+2$, or to $\mathbb{Q}$ for $s=0$. For conciseness, in all what follows, we will write $z=(x,y)$ and $z_{i}=(x_{i},y_{i})$ and $\tilde{z}_{j}=(\tilde{x}_{j},\tilde{y}_{j})$ for $1\leq i\leq n$ and $1\leq j\leq m$.
@@
$\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}\partial_{h}\delta_{\omega}^{out}\right]^{2}$
$\leq$ $\mathbb{E}_{\mathbb{Q}}\left[\sup_{\omega\in\Omega}(\partial_{h}\delta_{\omega}^{out})^{2}\right]$